### v1

# Frequency Decomposition-Driven Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

Abstract
Cross-domain semantic segmentation of remote sensing (RS) imagery based on unsupervised domain adaptation (UDA) techniques has significantly advanced deep-learning applications in the geosciences. Recently, with its ingenious and versatile architecture, the Transformer model has been successfully applied in RS-UDA tasks. However, existing UDA methods mainly focus on domain alignment in the high-level feature space. It is still challenging to retain cross-domain local spatial details and global contextual semantics simultaneously, which is crucial for the RS image semantic segmentation task. To address these problems, we propose novel high/low-frequency decomposition (HLFD) techniques to guide representation alignment in cross-domain semantic segmentation. Specifically, HLFD attempts to decompose the feature maps into high- and low-frequency components before performing the domain alignment in the corresponding subspaces. Secondly, to further facilitate the alignment of decomposed features, we propose a fully global-local generative adversarial network, namely GLGAN, to learn domain-invariant detailed and semantic features across domains by leveraging global-local transformer blocks (GLTBs). By integrating HLFD techniques and the GLGAN, a novel UDA framework called FD-GLGAN is developed to improve the cross-domain transferability and generalization capability of semantic segmentation models. Extensive experiments on two fine-resolution benchmark datasets, namely ISPRS Potsdam and ISPRS Vaihingen, highlight the effectiveness and superiority of the proposed approach as compared to the state-of-the-art UDA methods. The source code for this work will be accessible at https://github.com/sstary/SSRS.
keywords:
Remote Sensing, Semantic Segmentation, Unsupervised Domain Adaptation, Global-Local Information, High/Low-Frequency Decomposition
[…]
1 Introduction
Driven by the rapid development of earth observation technology, semantic segmentation has become one of the most important tasks in practical remote sensing (RS) applications, such as land cover monitoring, land planning, and natural disaster assessment. In the era of deep learning, semantic segmentation models are normally trained by exploiting labeled datasets in a fully supervised manner (Wang et al., 2022b; Zhang et al., 2022b). However, their impressive performance is limited by the availability of high-quality annotated samples generated by the costly and laborious labeling process. Furthermore, they show poor generalization when the trained models are applied to data in other domains or unseen scenarios due to discrepancies between heterogeneous domains. To circumvent the domain shift, the Unsupervised Domain Adaption (UDA) approach has been investigated, which aims to transfer knowledge from the source and target domains and achieve the desired semantic segmentation performance in the target domain.
Figure 1: The main challenges of UDA semantic segmentation in remote sensing images. (a) Intra-domain: Ground objects exhibit varying scales and irregular boundaries. (b) Inter-domain: Larger-scale variations and more complicated boundaries are observed across domains.
Generally speaking, the core idea of UDA methods is to learn domain-invariant features across domains based on domain alignment, including discrepancy-based (Zhang et al., 2015), reconstruction-based (Ghifary et al., 2016) and adversarial-based (Ganin et al., 2016; Zhang et al., 2018; Chen et al., 2019) optimization principles. In particular, the adversarial-based methods were developed upon the generative adversarial network (GAN) that alternately trains a generator and a discriminator for domain alignment (Tsai et al., 2018; Vu et al., 2019). In semantic segmentation tasks, adversarial-based methods attempt to align two domains at image-level (Tasar et al., 2020; Li et al., 2022a), feature-level (Gao et al., 2023; Ma et al., 2023b) or output-level (Tsai et al., 2018; Vu et al., 2019; Ni et al., 2023). Since global context modeling is important for RS vision tasks, transformer (Vaswani et al., 2017), with its superiority by its ability to model long-range contextual information, has facilitated many GAN-based UDA methods by enhancing feature alignment (Li et al., 2022b; Zhang et al., 2022a; Ma et al., 2023b) or boosting the discriminative capability of the discriminator (He et al., 2022; Tu et al., 2022).
Notably, UDA on semantic segmentation of remote sensing images presents unique challenges. For example, the ground objects and their spatial relationships are complex in fine-resolution remote sensing images (Zheng et al., 2020). As shown in Fig. 1(a), ground objects such as buildings and trees vary widely in scale, even in the same location (intra-domain). Furthermore, the boundaries of ground objects are irregular, as illustrated in Fig. 1(b), making feature extraction very challenging. Finally, the spatial relationship between ground objects is more implicit. These challenges become more evident when images are acquired from different locations or domains (inter-domain), which incurs serious domain shift problems in UDA tasks. Therefore, we argue that both detailed local features and abstract global contexts are crucial for domain alignment in semantic segmentation. Although current methods perform well in aligning high-level features that mainly contain global contexts, the discrepancy in local details between the source and target domains is often neglected, leaving a gap in knowledge transfer across domains. Furthermore, it is still challenging to learn multiscale domain-invariant representations, which are beneficial for semantic segmentation tasks.
To address these problems, we propose a frequency decomposition-driven UDA method based on a global-local GAN model, namely FD-GLGAN, considering alignment in low-frequency global representations and high-frequency local information. Firstly, we propose a high/low-frequency decomposition (HLFD) technology that decomposes the feature map into high-frequency and low-frequency components through a CNN-based extractor and a window-based attention module, respectively. By aligning high- and low-frequency information simultaneously, the generator can learn more transferable features from the source domain. In particular, the plug-and-play HLFD can be easily embedded into other UDA backbones as a universal module to facilitate the learning of domain-invariant representations. Considering that the HLFD aligns the high- and low-frequency information at the encoding stage, to further exploit the decomposed information in adversarial learning strategy, a novel GAN model named GLGAN is proposed by exploiting an HLFD-based generator and a global-local discriminator (GLDis) both of which are derived from the global-local transformer block (GLTB) (Wang et al., 2022b). With the exploration of global contexts and local spatial information, the former facilitates the generation of accurate segmentation maps whereas the latter efficiently distinguishes the origins of segmentation maps. The contributions of this work are summarized in the following:
  • A novel plug-and-play HLFD module is proposed to learn high- and low-frequency domain-invariant features at multiple scales based on frequency-based feature decomposition. To the best of our knowledge, this is the first work that introduces the idea of frequency decomposition into UDA for remote sensing image semantic segmentation.
  • We develop a fully global-local adversarial learning-based UDA framework. It facilitates domain alignment by capturing cross-domain dependency relationships at different levels and exploiting global-local context modeling between two domains.
  • Extensive experiments on two benchmark datasets i.e., ISPRS Potsdam and Vaihingen, confirm that the proposed FD-GLGAN is more effective in overcoming the domain shift problem in cross-domain semantic segmentation than existing UDA methods.
The remaining parts of this paper are organized as follows. We first review the related works in Sec. 2 and the related theories in Sec. 3. Sec. 4 introduces the proposed HLFD and GLGAN in detail, whereas Sec. 5 elaborates on the experiment setup, experimental results and related analysis. Finally, the conclusion is given in Sec. 6.
2 Related Work
2.1 Cross-Domain Semantic Segmentation in Remote Sensing
Leveraging encoder-decoder-based networks (Ronneberger et al., 2015), numerous methods of semantic segmentation have emerged in the field of remote sensing (Kemker et al., 2018; Kotaridis and Lazaridou, 2021). Most are developed based on CNNs with residual learning and attention mechanisms to boost the representation learning (Diakogiannis et al., 2020; Kotaridis and Lazaridou, 2021; Li et al., 2021).To address the limitations of CNNs in modeling global representations, many transformer-based methods (Wang et al., 2022a; Ma et al., 2024) have been integrated under the encoder-decoder framework, leveraging the global context modeling capability of the self-attention mechanism (Vaswani et al., 2017; Liu et al., 2021).
In contrast to the straightforward semantic segmentation task, cross-domain semantic segmentation focuses on transferring knowledge from the source domain to the target domain. According to the transfer strategy, existing cross-domain semantic segmentation models can be divided into two main categories, namely the self-training approach (Yan et al., 2019; Liang et al., 2023) and the adversarial training approach (Tsai et al., 2018; Vu et al., 2019). The former generates pseudo-labels for the target domain using a model trained on the labeled source domain by exploiting pseudo-labels for the target domain. Alternatively, the latter approach focuses on producing semantic segmentation maps for the target domain that resemble those in the source domain. This is achieved by aligning features or outputs, which can be facilitated through the use of Generative Adversarial Networks (GANs).
Table 1: GAN-based Model classification according to the utilization strategies of global and local information.
Generator Discriminator Method
Local
Local
GANAI (Benjdira et al., 2019), TriADA (Yan et al., 2019), CAGAN (Liu et al., 2022), CCAGAN (Chen et al., 2022), MASNet (Zhu et al., 2023), MBATA-GAN (Ma et al., 2023b)
Global
Local
MuGCDA (Xi et al., 2023), ALE-UDA (Zhang et al., 2022a)
Global
Global
TransGAN (Jiang et al., 2021), ViTGAN (Lee et al., 2021), Swin-GAN (Wang et al., 2022d), HyperViTGAN (He et al., 2022), SWCGAN (Tu et al., 2022)
Global-Local
Global-Local
GLGAN (Ours)
2.2 GAN-based UDA in Remote Sensing
GAN-based UDA algorithms have been proven effective in cross-domain semantic segmentation for remote sensing imagery. Benjdira et al. (2019) proposed the first GAN-based model designed for cross-domain semantic segmentation from aerial imagery. On this basis, TriADA (Yan et al., 2019) introduced a triplet branch to improve the CNN-based feature extraction by simultaneously aligning features from the source and target domains. Recently, a self-training approach that employs high-confidence pseudo-labels was applied to incorporate category information (Liang et al., 2023; Ma et al., 2023a; Shen et al., 2024). Furthermore, SADA (Song et al., 2019) and CS-DDA (Zhang et al., 2020) developed subspace alignment methods to constitute a shared subspace with high-level features. To improve the high-level feature extraction, attention mechanisms were introduced into GAN-based UDA. More specifically, CAGAN (Liu et al., 2022) developed a covariance-based channel attention module to compute weighted feature maps, whereas CCAGAN (Chen et al., 2022) used a category-certainty attention module to align the category-level features. MASNet (Zhu et al., 2023) stored domain-invariant prototypical representations by employing a feature memory module while MBATA-GAN (Ma et al., 2023b) aligned high-level features by constructing a cross-attention-based transformer (Xu et al., 2021; Wang et al., 2022e). Compared with the methods above, our work attempts to fully explore global-local context modeling in the generator and discriminator, as shown in Table 1. Therefore, it can capture cross-domain global semantics and local detailed contexts while boosting the learning of domain-invariant features.
2.3 Decomposition-based UDA in Remote Sensing
The core idea of decomposition-based UDA is to split the features into different components before performing domain alignment in the corresponding subspaces. For instance, ToMF-B (Zhao et al., 2022) decomposed the cross-domain features into task-related features and task-irrelevant features by analyzing the gradients of the predicted score corresponding to the labels. Furthermore, ST-DASegNet (Zhao et al., 2024) introduced a domain disentangled module to extract cross-domain universal features and purify single-domain distinct features in a self-training framework. However, these techniques mainly concentrate on aligning high-level abstract features, overlooking the relationship between high/low-frequency details and the characteristics of ground objects. Therefore, it is crucial to address this shortfall by delving into frequency decomposition technology.
3 Preliminaries
Before introducing our proposed framework for UDA, we first revisit the basics of the multi-head self-attention (MHSA) design and the convolutional operation (Conv) from a frequency perspective. We employ the discrete Fourier transform (DFT) to analyze the feature maps regarding specific frequency components generated by MHSA and Conv. Let $X\in\mathbb{R}^{H\times W}$ and $Y\in\mathbb{R}^{C}$ represent an image in the spatial domain and its corresponding label, where $C$ denotes the number of classes. We transform $X$ into the frequency spectrum using the DFT $\mathcal{F}$: $\mathbb{R}^{H\times W}\rightarrow\mathbb{C}^{H\times W}$ and revert the signals of the image from frequency back to the spatial domain using the inverse DFT $\mathcal{F}^{-1}$: $\mathbb{C}^{H\times W}\rightarrow\mathbb{R}^{H\times W}$. For a mask $m\in\{0,1\}^{H\times W}$, the low-pass filtering $\mathcal{M}^{S}_{\mathrm{low}}$ and high-pass filtering $\mathcal{M}^{S}_{\mathrm{high}}$ with the filter size $S$ are formally defined as (Bai et al., 2022):
$\mathcal{M}^{S}_{\mathrm{low}}=\mathcal{F}^{-1}(m\odot\mathcal{F}(x)),\mathrm{where~{}}m_{i,j}=\begin{cases}1,~{}\mathrm{if~{}min}(|i-\frac{H}{2}|,|j-\frac{W}{2}|)\leq\frac{S}{2}\\ 0,~{}\mathrm{otherwise}\end{cases},$ (1)
and
$\mathcal{M}^{S}_{\mathrm{high}}=\mathcal{F}^{-1}(m\odot\mathcal{F}(x)),\mathrm{where~{}}m_{i,j}=\begin{cases}0,~{}\mathrm{if~{}min}(|i-\frac{H}{2}|,|j-\frac{W}{2}|)\leq\frac{\mathrm{min}(H,W)-S}{2}\\ 1,~{}\mathrm{otherwise}\end{cases},$ (2)
where $\odot$ stands for the Hadamard product, and $m_{i,j}$ denotes the value of $m$ at position $(i,j)$. According to the definitions in Eq. (1) and Eq. (2), the low-frequency components are shifted to the center of the frequency spectrum while the high-frequency components are shifted away from the center of the frequency spectrum.
Figure 2: The basic ideas of our frequency decomposition strategy. (a) Feature map extracted by the encoder, (b) feature map extracted further by MHSA and (c) its corresponding frequency spectrum; (d) feature map extracted further by Conv and (e) its corresponding frequency spectrum. The centroid represents low-frequency information, while the distance from it represents high-frequency information. More high-value points represent more information for the corresponding frequency.
Fig. 2 illustrates the contrasting behaviors that MHSA and Conv demonstrate. As shown in Fig. 2(b) and (d), MHSA aggregates feature maps, whereas Conv disperses them. Furthermore, as depicted in Fig. 2(c) and (e), Fourier analysis of feature maps reveals that MHSA focuses on low-frequency components, while Conv amplifies high-frequency components. In other words, MHSAs function as low-pass filters, while Conv serves as high-pass filters (Park and Kim, 2022). Consequently, MHSA and Conv complement each other and can be jointly employed for frequency decomposition. More details can be found in (Wang et al., 2020, 2022c).
Figure 3: The overview of the proposed FD-GLGAN that is comprised of a Generator, a Discriminator and the proposed HLFDs.
4 Proposed Method
In this section, we will first propose the overall structure of FD-GLGAN before elaborating on the proposed HLFD module as well as its generator and discriminator (GLDis). After that, the loss functions used in training will be presented.
Figure 4: The detailed structure of encoder and the detailed structure of the SwinTB.
4.1 Framework
Typical UDA tasks involve a labeled source dataset $D_{s}=\left\{(X_{s},Y_{s})\right\}^{n_{s}}$ and an unlabeled target dataset $D_{t}=\left\{X_{t}\right\}^{n_{t}}$ where $X_{s}$ represents an image in the source domain with its corresponding label $Y_{s}$ whereas $X_{t}$ denotes the unlabeled image in the target domain. Furthermore, $n_{s}$ and $n_{t}$ stand for the sample size of the source domain and the target domain, respectively. As the source domain and target domain possess different marginal and conditional distributions, i.e. $p_{s}(X_{s})\neq p_{t}(X_{t})$, the domain shift problem occurs. We assume that certain inherent similarities exist between different domains, which is valid as both domains share the same segmentation output space (Tsai et al., 2018; Wang et al., 2022e; Ma et al., 2023b).
As depicted in Fig. 3, the proposed FD-GLGAN is designed by capitalizing upon frequency decomposition and global-local context utilization. Specifically, a novel HLFD module is presented to align cross-domain representations by exploiting multiscale features generated by the encoder. Moreover, a novel global-local discriminator named GLDis further exploits the decomposed features to learn multiscale domain-invariant representations. In particular, the GLDis attempts to mimic the decoder to learn the domain-specific output space using the global-local representations.
4.2 Encoder
As illustrated in Fig. 4, the encoder comprises a Patch Partition layer and four successive stages. More specifically, the first stage contains one Embedding layer followed by the SwinTBs, whereas the following three stages are equipped with one Merging layer followed by SwinTBs each. We denote by $X_{s}\in\mathbb{R}^{{H\times W\times Z}}$ and $X_{t}\in\mathbb{R}^{{H\times W\times Z}}$ the training images from the source and target domains, respectively. Furthermore, $Z$ stands for the number of image channels, whereas $H$ and $W$ represent the height and width of the image, respectively. $X_{s}$ and $X_{t}$ are first split into non-overlapping patches of size $4\times 4$ by the Patch Partition layer. After that, the first-stage Embedding layer is applied to project these patches onto the encoding space before the SwinTBs gather image information on these patch tokens. As shown in Fig. 4, the SwinTB consists of four LayerNorm (LN) layers, two multilayer perceptrons (MLPs), one window-based multi-head self-attention (W-MHSA) and one shifted window-based multi-head self-attention (SW-MHSA) (Vaswani et al., 2017; Liu et al., 2021). After the first stage, three more successive stages are applied to generate the multiscale features denoted as $F_{s}^{\emph{1-4}}$ and $F_{t}^{\emph{1-4}}$ both of size ${\mathcal{S}}_{i}\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}$ where ${\mathcal{S}}_{i}$ is the encoding dimension with $i\in\left\{1,2,3,4\right\}$ being the stage index. After that, these multiscale features are then fed into the corresponding HLFD module for frequency decomposition. Notably, the images from the source and the target domain are fed into the generator successively, that is, $F_{s}^{\emph{1-4}}$ and $F_{t}^{\emph{1-4}}$ are generated and collected, respectively.
Figure 5: The dataflow of the proposed multiscale HLFDs. The multiscale features from the source and target domains are decomposed by the corresponding scale HLFD into low-frequency and high-frequency decomposition features.
[…]
4.3 HLFD module
HLFD modules are employed to decompose the multiscale feature maps $F_{s}^{{1-4}}$ and $F_{t}^{{1-4}}$, as illustrated in Fig. 5. The feature map at each scale, including the features from the source domain and the target domain, will be decomposed into the high or low-frequency feature maps by the corresponding scale HLFD before these outputs are used to calculate the frequency alignment loss. Fig. 6 depicts the structure of the proposed HLFD module. Specifically, the HLFD module comprises two parallel branches, namely the low-frequency and high-frequency extractor. The former captures the low-frequency components by utilization of the W-MHSA (Liu et al., 2021). The following cross-shaped window context interaction (WCI) can maintain the global spatial consistency of ground objects. In contrast, the latter, designed for high-frequency components, has a structure of two parallel convolutional layers with kernel sizes of $3$ and $1$, respectively. Each convolutional layer is followed by a batch normalization operation. These branches facilitate the frequency decomposition of multiscale features based on the discussion in Sec. 3. The output low- and high-frequency components are then processed by an LN layer and an MLP, respectively. The final outputs of HLFD module denoted as $*LF_{s}^{\emph{*}}/LF_{t}^{\emph{*}}$ and $*HF_{s}^{\emph{*}}/HF_{t}^{\emph{*}}$ are used to computed the frequency alignment loss $L_{\rm{f}}$ that will be elaborated in Sec. 4.6.
Figure 7: (a) The detailed structure of GLTB. (b) The detailed structure of the decoder. (c) The detailed structure of the discriminator
4.4 Decoder
The decoder and GLDis that will be presented in Sec. 4.5 are both based on GLTB and designed to further exploit the decomposed features by gradually extracting global and local information. Their detailed structures and the GLTB are depicted in Fig. 7. Comparison of Fig. 6 and Fig. 7(a) suggests that the GLTB and HLFD modules share structural similarities. Specifically, the HLFD is replaced by global-local attention constituted of two parallel branches, namely the global and the local branches. The architecture of these branches closely mirrors that of the frequency extractor. However, in contrast to HLFD that is utilized for feature decomposition, the global-local attention leverages its two branches to investigate and incorporate global-local information. Hereby, the outputs from both the global and local branches are fused together before the resulting global–local context is further characterized by two convolution layers and a batch normalization operation.
The high-level feature maps derived from the encoder are first fed into an individual GLTB whose output is processed by two stages of weighted sum operation and GLTB. In particular, GLTB is able to capture the global context and local details simultaneously. After that, the weighted sum operation is employed to extract domain-invariant features by adaptively fusing the residual features and those from the previous GLTB, as shown in Fig. 7(b). After the first three stages of the decoder, a weighted sum operation is used to fuse the residual connection from the first stage of the encoder and the deep global-local feature derived from previous GLTB before a classifier produces the final segmentation maps denoted by $P_{s}$ and $P_{t}$ of size $\mathbb{R}^{H\times W\times C}$ for the source and target domains, respectively.
Finally, the supervised segmentation loss $L_{\rm{seg}}$ and the adversarial loss $L_{\rm{adv}}$ are also derived in the generator optimization stage. More specifically, $L_{\rm{seg}}$ is computed based on the source labels while $L_{\rm{adv}}$ is calculated using the adversarial strategy. More details about the loss functions will be elaborated in Sec. 4.6.
4.5 Discriminator
Inspired by UNetformer (Wang et al., 2022b), we employ GLTB to exploit the semantic output space produced by the generator. As shown in Fig. 7(c), the proposed GLDis adopts a structure similar to the decoder but with inverted sampling operations. Specifically, it first down-samples the weighted self-information (WSI) distribution $I_{s}$ and $I_{t}$ derived from $P_{s}$ and $P_{t}$ before feeding the down-sampled data into a GLTB. More specifically, given the output segmentation maps of the source domain ${P}_{s}$ and the target domain ${P}_{t}$, their self-information are defined as $-\mathrm{log}{P}_{s}^{(H,W,c)}$ and $-\mathrm{log}{P}_{t}^{(H,W,c)}$, thereby the WSI maps ${I}_{s}$ and ${I}_{t}$ are computed by:
${I}_{s}^{(H,W,c)}=-{P}_{s}^{(H,W,c)}\cdot\mathrm{log}{P}_{s}^{(H,W,c)},$ (3)
${I}_{t}^{(H,W,c)}=-{P}_{t}^{(H,W,c)}\cdot\mathrm{log}{P}_{t}^{(H,W,c)}.$ (4)
These information maps are regarded as the disentanglement of the Shannon Entropy. After the first GLTB, the resulting data is further processed by two successive stages of down-sampling followed by the GLTB, designed to gradually gather the global and local information in the segmentation map-generated features. Finally, the resulting data is down-sampled before being processed by a Classifier, which gives the predictions of the domain to which the input segmentation maps belong. Despite its appearance similar to the decoder structure, GLDis is tailored to learn domain-specific representations using an adversarial strategy. More specifically, the decoder upsamples the abstract semantic features, whereas GLDis down-samples the structured output space to identify the domain to which each input belongs. The final output of GLDis is a cross-entropy loss denoted by $L_{\rm{d}}$ in the discriminator training stage. More details about the loss function will be presented in Sec. 4.6.
4.6 Loss Functions
The proposed FD-GLGAN capitalizes on three loss functions, namely the segmentation loss $L_{\rm{seg}}$, the typical adversarial loss $L_{\rm{adv}}$, the frequency alignment loss $L_{\rm{f}}$, and the cross-entropy loss $L_{\rm{d}}$. These loss functions are defined as follows.
4.6.1 Generator
$L_{\rm{seg}}$, $L_{\rm{f}}$ and $L_{\rm{adv}}$ are proposed to optimize the generator. In particular, $L_{\rm{seg}}$ is computed directly by images from the source domain and their corresponding labels. Mathematically, $L_{\rm{seg}}$ takes the following form:
$L_{\rm{seg}}({P}_{s})=-\lambda^{seg}\sum_{H,W}\sum_{c\in{\mathcal{C}}}Y_{s}^{(H,W,c)}\left[\mathrm{log}({P}_{s}^{(H,W,c)})\right],$ (5)
where ${\mathcal{C}}$ is the category set whose cardinality is $C$ while $Y_{s}$ and $\lambda^{seg}$ denote the labels from the source domain and the weighting coefficient for the segmentation loss, respectively.
$L_{\rm{f}}$ is computed to align decomposed cross-domain features generated by multiscale HLFDs as follows:
$\begin{split}&L_{\rm{f}}\left(LF_{s}^{\emph{1-4}},LF_{t}^{\emph{1-4}},HF_{s}^{\emph{1-4}},HF_{t}^{\emph{1-4}}\right)\\ &=L_{\rm{l}}\left(LF_{s}^{\emph{1-4}},LF_{t}^{\emph{1-4}}\right)+L_{\rm{h}}\left(HF_{s}^{\emph{1-4}},HF_{t}^{\emph{1-4}}\right)\\ &=\lambda^{f}\cdot\displaystyle\sum_{i=1}^{\emph{4}}\displaystyle\sum_{d=1}^{D_{i}}\frac{\left(\mathcal{P}(LF_{s}^{i,d})-\mathcal{P}(LF_{t}^{i,d})\right)}{D_{i}}+\lambda^{f}\cdot\displaystyle\sum_{i=1}^{\emph{4}}\displaystyle\sum_{d=1}^{D_{i}}\frac{\left(\mathcal{P}(HF_{s}^{i,d})-\mathcal{P}(HF_{t}^{i,d})\right)}{D_{i}},\end{split}$ (6)
where $\lambda^{f}$ is a weighting coefficient, $D_{i}$ is the number of channels for the $i$-th scale, and $\mathcal{P}$ stands for the average pool, which gathers spatial information for each channel $d$. The frequency alignment loss function is designed to efficiently align the feature maps across the source and target domains.
To align the structured output space, the WSI distribution is adopted in the proposed training process to calculate the adversarial loss as follows:
$L_{\rm{adv}}\left({I}_{t}\right)=-\lambda^{adv}\sum_{H,W}\left[\mathrm{log}\left(\mathcal{D}({I}_{t})^{(H,W,0)}\right)\right],$ (7)
where $\lambda^{adv}$ is the weighting coefficient for the adversarial loss while $\mathcal{D}$ denotes the GLDis. Note that the artificial labels 00 and $1$ are assigned to samples drawn from the source domain and the target, respectively. In Eq. (7), the label 00 in the superscript ${}^{(H,W,0)}$ is set to intentionally mislead GLDis. Therefore, the discriminator is motivated to bridge the prediction gap between the source and target domains.
Finally, the following overall objective function is proposed to optimize the generator:
$L_{G}=L_{\rm{seg}}+L_{\rm{f}}+L_{\rm{adv}}.$ (8)
4.6.2 GLDis
The discriminator GLDis is optimized with the following cross-entropy loss $L_{\rm{d}}$ to identify to which domain a given segmentation map belongs.
$L_{\rm{d}}({I}_{s},{I}_{t})=-\sum_{H,W}\left[\mathrm{log}(\mathcal{D}({I}_{s})^{(H,W,0)})+\mathrm{log}(\mathcal{D}({I}_{t})^{(H,W,1)})\right].$ (9)
4.6.3 Overall Loss Function
Finally, the overall loss function proposed for FD-GLGAN is given as follows:
$\mathop{\max}_{\mathcal{D}}\ \mathop{\min}\limits_{\mathcal{G}}L(X_{s},Y_{s},X_{t}),$ (10)
where $\mathcal{D}$ and $\mathcal{G}$ denote GLDis and the generator, respectively. Eq. (10) is designed to minimize the segmentation loss in the generator for the source images while maximizing the output space similarity of the source and target domains. Notably, the generator and the discriminator are trained in a classical GAN-based manner, whereas the generator and GLDis are trained alternatively. More specifically, the generator is first trained using Eq. (8) with the parameters of GLDis being held constant before GLDis is updated with Eq. (9).
5 Experiments And Discussions
5.1 Datasets Description
To verify the effectiveness and robustness of the proposed FD-GLGAN, extensive experiments were conducted on two open-source benchmark datasets (Gerke, 2014) involving fine-resolution images obtained from two cities (Potsdam [^1] https://www.isprs.org/education/benchmarks/UrbanSemLab/2d-sem-label-potsdam.aspx and Vaihingen [^2] https://www.isprs.org/education/benchmarks/UrbanSemLab/2d-sem-label-vaihingen.aspx). The images in both datasets were obtained from aerial photography but of different Ground Sampling Distance (GSD). More specifically, the images of Potsdam have clearer ground objects with a GSD of $5$ cm as compared to a GSD of $9$ cm for Vaihingen. Furthermore, images in Potsdam possess richer spectrum information from four frequency bands, namely, InfraRed, Red, Green and Blue (IRRGB), whereas images in Vaihingen only three bands, namely, InfraRed, Red and Green channels (IRRG). In addition, the difference in geographical locations results in more prominent discrepancies even for the same ground object categories due to different road layouts, building types and vegetation forms. These discrepancies present a major challenge to UDA methods.
Figure 8: Image samples of size $1024\times 1024$ from (a) P-RGB, (b) P-IRRG and (c) V-IRRG, respectively. Despite some visual similarities between these two datasets, they exhibit large discrepancies in the distribution of land objects.
1) Potsdam: There are $24$ very fine-resolution true orthophotos of size $6000\times 6000$ in Potsdam dataset. In the experiments, we divided these orthophotos into a training set of $18$ patches and a test set of $6$ patches as follows:
  • The training set (18 patches): {6_10, 7_10, 2_12, 3_11, 2_10, 7_8, 5_10, 3_12, 5_12, 7_11, 7_9, 6_9, 7_7, 4_12, 6_8, 6_12, 6_7, 4_11};
  • The test set (6 patches): {2_11, 3_10, 4_10, 5_11, 6_11, 7_12}.
2) Vaihingen: The Vaihingen dataset contains $16$ very fine-resolution true orthophotos of size $2500\times 2000$ pixels. In our experiments, these orthophotos are divided into a training set of $12$ patches and a test set of $4$ patches as follows:
  • The training set (12 patches): {1, 3, 23, 26, 7, 11, 13, 28, 17, 32, 34, 37};
  • The test set (4 patches): {5, 21, 15, 30}.
All images in both datasets are provided with semantic labels, including six ground object classes, namely, Building (Bui.), Tree (Tre.), Low Vegetation (Low.), Car, Impervious Surface (Imp.) and Clutter/Background. Fig. 5 illustrates two groups of images sampled from the Potsdam and Vaihingen datasets. Visual inspection of Fig. 5 confirms large discrepancies across the two datasets. In the sequel, we divided the Potsdam dataset into two subsets, namely, “P-IRRG” and “P-RGB”, taking into account the data size and image channels as compared to those of Vaihingen. The Vaihingen dataset forms only one dataset denoted as “V-IRRG”. By treating these three subsets as the source and target domains, we designed three UDA tasks with increasing levels of difficulty as listed in Table 2. It is worth noting that these three tasks exhibit different levels of domain shifts. For instance, Task $1$, i.e. P-IRRG to V-IRRG, suffers from the domain shift caused by large discrepancies in the spatial resolution, the ground object style and the illumination environment, though the images from both domains possess the same spectrum information. In contrast, the domain shift problem in Task $2$, i.e. P-RGB to V-IRRG, is caused by the frequency band difference in addition to all discrepancies in Task $1$. Finally, Task $3$, i.e. V-IRRG to P-RGB, encounters the same domain shift problem as Task $2$ with a greater challenge as the size of the Vaihingen dataset in the source domain is much smaller than that of the Potsdam dataset in the target domain. In summary, these three UDA tasks were designed to explore the effectiveness of the UDA method under consideration in cross-location, in-image mode, and cross-image mode UDA tasks.
Table 2: Unsupervised Domain Adaptation Tasks
Source Domain Target Domain
Task 1 P-IRRG V-IRRG
Task 2 P-RGB V-IRRG
Task 3 V-IRRG P-RGB
Table 3: Nine methods were evaluated in our following experiments, including two non-UDA and seven UDA methods.
Index Method Reference
Description
1 Source-only (Jha et al., 2019)
Images and labels from the source domain are utilized to train the segmentation network before it is tested directly on the target domain.
2 DANN (Ganin et al., 2016)
The first UDA method that applies adversarial learning to solve the domain shift problem on the image classification task.
3 AdasegNet (Tsai et al., 2018)
This UDA method performs multi-level adversarial learning with two discriminators while assuming that the source and target domains share a similar output space.
4 Advent (Vu et al., 2019)
This UDA method aligns the entropy maps of the output derived by the generator for the source and target domains.
5 GANAI (Benjdira et al., 2019)
The first UDA method designed for semantic segmentation of remote sensing images based on the GAN framework.
6 TriADA (Yan et al., 2019)
This UDA method employs a triplet branch, considering both domains simultaneously to learn a domain-invariant classifier using a domain similarity discriminator.
7 CCAGAN (Chen et al., 2022)
This UDA method develops a category-certainty attention module focusing on the misaligned regions in the category level.
8 MBATA-GAN (Ma et al., 2023b)
This UDA method develops a mutually boosted attention-based transformer to augment the high-level features and a feature discriminator to learn the transferable features.
9 Supervised (Jha et al., 2019)
Images and labels from the target domain are used to train and test the segmentation network, which stands for the optimal performance that any UDA method can expect.
5.2 Evaluation Metrics
In this work, we adopt three metrics to quantitatively assess the segmentation performance of the cross-domain remote sensing images for all UDA methods, namely, the Overall Accuracy (OA), the F1 score (F1) and the Intersection over Union (IoU), which are defined as follows:
$OA$ $=$ $\frac{TP+TN}{TP+TN+FP+FN},$ (11)
$Precision$ $=$ $\frac{TP}{TP+FP},$ (12)
$Recall$ $=$ $\frac{TP}{TP+FN},$ (13)
$F1$ $=$ $2\times\frac{Precision\times Recall}{Precision+Recall},$ (14)
$IoU$ $=$ $\frac{TP}{TP+FP+FN},$ (15)
where $TP$, $FP$, $TN$ and $FN$ are true positives, false positives, true negatives and false negatives, respectively. Furthermore, the mean F1 score (mF1) and the mean IoU (mIoU) for the main five classes apart from Clutter/Background are calculated to measure the average performance of the UDA methods following the common practice reported in the literature (Yan et al., 2019; Ma et al., 2023b; Yan et al., 2021).
In addition, the computational complexity of the proposed GLGAN and FD-GLGAN is evaluated using the following metrics. First, the floating point operation count (FLOPs) is employed to evaluate the model complexity. Second, the number of model parameters (M) and the memory footprint (MB) are utilized to evaluate the memory requirement. Finally, the frames per second (FPS) is used to evaluate the execution speed. For a computationally efficient method, its first three metrics are usually small, while its FPS value should be large.
For comparison purposes, we benchmark the proposed model against several state-of-the-art models trained for $50$ epochs, each containing $1000$ batches. All models are evaluated every $100$ batch before their performances are recorded during the training stage. Finally, each method’s best model corresponding to the highest segmentation performance is recorded.
Table 4: Task 1: Experimental results under the variations in geographic locations only: Quantitative comparison (%) with the state-of-the-art on the adaptation P-IRRG to V-IRRG. We show F1 for each category and the best values are bolded except Supervised method.
Method Backbone Bui. Tre. Low. Car Imp. OA mF1 mIoU
Non-DA (Source-Only)
ResUNet++
76.93
70.79
40.19
7.06
56.70
63.03
50.33
37.13
DANN
83.07
75.93
54.15
55.38
69.73
71.82
67.66
52.24
AdasegNet
ResNet101
87.87
79.44
56.73
57.59
76.10
76.29
71.55
57.14
Advent
ResNet101
90.85
78.78
58.31
55.07
79.53
77.63
72.51
58.68
GANAI
ResNet101
87.13
77.20
52.48
58.78
74.38
73.52
69.99
55.29
TriADA
ResNet101
89.34
79.06
47.50
65.72
81.11
77.56
72.55
58.88
CCAGAN
ResNet101
89.02
78.10
59.29
65.42
79.10
76.90
74.19
60.09
MBATA-GAN
ResNet101
90.98
79.51
63.41
66.38
82.80
80.75
76.77
63.50
GLGAN
Swin-Base
91.95
80.31
67.71
65.89
84.18
81.97
78.01
65.04
Swin-Large
92.33
81.60
69.11
67.34
84.58
82.74
78.99
66.30
FD-GLGAN
Swin-Base
93.94
80.95
70.06 70.40 86.13 83.66 80.30 68.09
Non-DA (Supervised)
ResUNet++
94.84
90.15
78.25
80.16
90.47
89.66
86.77
77.20
5.3 Experimental settings
In our work, UNetformer (Wang et al., 2022b) was adopted as the generator, which integrated the Swin Transformer (Liu et al., 2021) pretrained on ImageNet (Deng et al., 2009) as its encoder. In our experiments, two versions of Swin Transformers, namely Swin-Base and Swin-Large, were applied for image feature extraction. Both Swin-Base and Swin-Large have four stages with the stack of $\left\{2,2,18,2\right\}$ SwinTBs with the embedding dimensions of Swin-Base and Swin-Large being $\left\{128,256,512,1024\right\}$ and $\left\{196,384,768,1536\right\}$, respectively. Furthermore, the number of the heads in Swin-Base and Swin-Large are $\left\{4,8,16,32\right\}$ and $\left\{6,12,24,48\right\}$, respectively. In addition, the window size for W-MHSA and SW-MHSA is $16\times 16$. For both decoder and GLDis, their decoder dimension is $256$, whereas the window size of the global-local attention is $8\times 8$.
All models in the experiments were implemented with the PyTorch framework on a single NVIDIA GeForce RTX 3090 GPU with $24$-GB RAM. To efficiently read the large RS patches, the sliding window method was adopted to collect the training samples dynamically. During the training stage, the size of the sliding window was set to $256\times 256$ with no overlapping, whereas the batch size was $10$. Furthermore, the generator was optimized with the Stochastic Gradient Descent (SGD) optimizer and the Nesterov acceleration with a momentum of $0.9$ and a weight decay of $5\times 10^{-4}$. The initial learning rate was set to $2.5\times 10^{-4}$, decreasing with a polynomial decay of power $0.9$. For GLDis, the Adam optimizer was deployed with the learning rate of $10^{-4}$ with a polynomial decay of power $0.9$. The momentum of Adam is set to $\left[0.9,0.99\right]$. Finally, the multi-level adversarial learning scheme was also adopted with parameters set to their default values derived from AdasegNet (Tsai et al., 2018) in which $\lambda^{seg}$ and $\lambda^{adv}$ were set to $1.0$ and $0.001$, respectively. $\lambda^{f}$ was set to $0.01$ according to the parameter analysis that will be presented in Sec. 5.7.
Figure 9: Task 1: qualitative performance comparison of the adaptation P-IRRG to V-IRRG with the size of $2048\times 2048$. (a) RGB images, (b) Ground Truth, (c) Source-only, (d) AdasegNet, (e) Advent, (f) GANAI, (g) TriADA, (h) CCAGAN, (i) MBATA-GAN, (j) GLGAN, (k) FD-GLGAN and (l) Supervised method.
5.4 Benchmarking
To confirm the effectiveness of the proposed methods, we performed comparison experiments on two non-DA methods, namely a deep model without domain adaptation (labeled as “Source-only”) and a supervised deep model (labeled as “Supervised”) trained with labeled data from the target domain. These two non-DA methods stand for the two extreme cases, with the “Supervised” model representing the optimal performance that any UDA method targets and the “Source-only” model standing for the performance lower bound. Furthermore, the classical ResUNet++ (Jha et al., 2019) is selected for these two non-DA methods by taking into account the balance between segmentation performance and computational complexity. In addition to the two non-UDA methods, some representative existing UDA methods were also selected for quantitative comparison. Table 3 summarizes the nine methods evaluated in our following experiments.
For the aforementioned methods, the same network architectures, training parameters, and evaluation methods were applied to make a fair comparison.
5.5 Results and Discussions
In this section, extensive experimental results are reported for the three domain adaptation tasks listed in Table 2.
Table 5: Task 2: Experimental results under the variations in both geographic locations and imaging mode: Quantitative comparison (%) with the state-of-the-art on the adaptation P-RGB to V-IRRG. We show F1 for each category and the best values are bolded except Supervised method.
Method Backbone Bui. Tre. Low. Car Imp. OA mF1 mIoU
Non-DA (Source-Only)
ResUNet++
64.48
0.04
5.26
1.57
58.21
35.71
25.91
18.43
DANN
73.20
62.46
41.50
54.94
64.39
61.85
59.30
42.94
AdasegNet
ResNet101
74.04
75.16
40.85
57.02
62.12
64.75
61.84
45.92
Advent
ResNet101
83.29
77.85
42.37
49.92
65.73
67.26
63.83
48.84
GANAI
ResNet101
76.87
56.25
41.50
57.93
66.07
61.81
59.72
43.57
TriADA
ResNet101
77.86
73.85
36.57
58.70
66.48
66.09
62.69
47.20
CCAGAN
ResNet101
85.60
75.27
36.31
63.24
66.03
68.38
65.25
50.58
MBATA-GAN
ResNet101
87.67
75.30
39.06
64.95
67.33
69.29
66.86
52.31
GLGAN
Swin-Base
87.71
68.37
58.59
66.68
73.73
73.21
71.02
55.98
Swin-Large
89.70
77.53
58.93
63.21
78.48
77.23
73.57
59.44
FD-GLGAN
Swin-Base
92.42
77.53
62.88 69.38 80.42 79.16 76.53 63.09
Non-DA (Supervised)
ResUNet++
94.84
90.15
78.25
80.16
90.47
89.66
86.77
77.20
Figure 10: Task 2: qualitative results of the adaptation P-RGB to V-IRRG with the size of $2048\times 2048$. (a) RGB images, (b) Ground Truth, (c) Source-only, (d) AdasegNet, (e) Advent, (f) GANAI, (g) TriADA, (h) CCAGAN, (i) MBATA-GAN, (j) GLGAN, (k) FD-GLGAN and (l) Supervised method.
Task 1 (P-IRRG to V-IRRG): The results for Task $1$ are summarized in Table 4. Inspection of Table 4 suggests that the “Source-Only” method showed the worst performance among all methods under consideration in terms of OA, mF1 and mIoU. Furthermore, all UDA methods demonstrated performance improvement of different degrees. For instance, the simplest DANN achieved improvements of $8.79\%$ in OA, $17.33\%$ in mF1 and $15.11\%$ in mIoU, which confirmed the necessity and effectiveness of the UDA approach. In particular, the baseline Advent showed improved OA, mF1, and mIoU of $77.63\%$, $72.51\%$ and $58.68\%$, respectively. Compared with the baseline, the proposed GLGAN achieved great improvements on all matrices. Specifically, the OA, mF1 and mIoU attained by GLGAN (Swin-Base) were $81.97\%$, $78.01\%$ and $65.04\%$, respectively, which amounts to an improvement of $4.34\%$, $5.50\%$ and $6.36\%$ as compared to Advent. In addition, GLGAN (Swin-Large) attained higher scores in OA, mF1 and mIoU of $82.74\%$, $78.99\%$ and $66.30\%$, respectively, by employing a more complex encoder. Among all the unsupervised methods examined, the proposed FD-GLGAN achieved the highest scores on all matrices with the smallest performance gap as compared to the “Supervised” non-DA method. It is easy to observe in Table 4 that the proposed FD-GLGAN attained the highest F1 score in four main categories, including $93.94\%$ for Buildings, $70.06\%$ for Low Vegetation, $70.40\%$ for Cars and $86.13\%$ for Impervious Surface, which amounts to improvement of $3.09\%$ for Buildings, $11.75\%$ for Low Vegetation, $15.33\%$ for Cars and $6.6\%$ for Impervious Surface. Upon comparing the results of GLGAN (Swin-Base) and FD-GLGAN (Swin-Base), it becomes evident that improvements are observed across all categories, suggesting that the HLFD module facilitates the learning of transferable features.
Finally, Fig. 9 presents the semantic segmentation results achieved by all methods under consideration. Visual inspection of Fig. 9 reveals that the proposed GLGAN and FD-GLGAN are more accurate in differentiating visually similar categories such as Trees and Low Vegetation. Furthermore, the segmentation boundary of ground objects provided by our method is much smoother, with fewer random prediction points than other UDA methods.
Task 2 (P-RGB to V-IRRG): Table 5 shows the results for Task $2$ that encounters a more severe domain shift problem than Task $1$. In addition to the discrepancies in the spatial resolution, the geographical location and other environmental factors, Task $2$ also has to learn image representations from the RGB bands before transferring the learned features to the IRRG bands. Notably, the “Source-only” method showed the worst performance, which is evidenced by its poor OA, mF1, and mIoU of $35.71\%$, $25.91\%$ and $18.43\%$, respectively. In sharp contrast, it is observed from Table 5 that the proposed GLGAN (Swin-Large) achieved OA of $77.23\%$, mF1 of $73.57\%$, and mIoU of $59.44\%$, which stands for an increase of $9.97\%$, $9.74\%$, and $10.60\%$ as compared to the corresponding performance of Advent, respectively. Further, the proposed FD-GLGAN (Swin-Base) achieved a higher performance with an OA of $79.16\%$, mF1 of $76.53\%$, and mIoU of $63.09\%$ than the GLGAN (Swin-Large). It proved the broad applicability of the proposed HLFD, which can steadily improve the performance of the original UDA method in different scenarios. Moreover, these results demonstrate that the proposed FD-GLGAN can effectively handle large domain shifts in Task $2$. These results demonstrate that the proposed GLGAN and FD-GLGAN can effectively handle large domain shifts in Task $2$.
Table 6: Task 3: Experimental results on domain shift in both geographic location and imaging mode: Quantitative comparison (%) with the state-of-the-art on the adaptation V-IRRG to P-RGB. We show F1 for each category and the best values are bolded except Supervised method.
Method Backbone Bui. Tre. Low. Car Imp. OA mF1 mIoU
Non-DA (Source-Only)
ResUNet++
42.87
0.10
7.36
3.55
43.27
34.14
19.43
12.11
DANN
63.70
42.50
40.26
58.21
62.10
54.53
53.35
37.00
AdasegNet
ResNet101
73.23
46.23
33.88
59.48
71.61
58.38
56.89
41.27
Advent
ResNet101
71.13
46.62
31.70
66.76
70.95
57.30
57.43
41.90
GANAI
ResNet101
64.97
38.79
45.29
67.45
67.17
57.02
56.73
40.58
TriADA
ResNet101
73.95
38.06
47.87
70.46
72.55
61.34
60.58
44.99
CCAGAN
ResNet101
76.45
44.86
47.28
72.82
71.26
61.68
62.53
46.87
MBATA-GAN
ResNet101
81.54
30.75
58.14
69.83
75.37
65.82
63.13
48.42
GLGAN
Swin-Base
82.77
60.46
65.19
77.27
76.14
71.15
72.37
57.34
Swin-Large
85.65
62.08
69.26
81.37
76.09
73.63
74.89
60.58
FD-GLGAN
Swin-Base
85.65 65.61 71.02
79.43
79.66 75.54 76.27 62.17
Non-DA (Supervised)
ResUNet++
92.08
74.91
80.13
90.79
89.35
84.69
85.46
75.19
Figure 11: Task 3: qualitative results of the adaptation V-IRRG to P-RGB with the size of $2048\times 2048$. (a) RGB images, (b) Ground Truth, (c) Source-only, (d) AdasegNet, (e) Advent, (f) GANAI, (g) TriADA, (h) CCAGAN, (i) MBATA-GAN, (j) GLGAN, (k) FD-GLGAN and (l) Supervised method.
Furthermore, it is interesting to observe that the FD-GLGAN improved the accuracy for Low Vegetation by $20.51\%$, but suffered from performance degradation of $0.32\%$ for Trees. This observation could be explained by the fact the trees in Potsdam are mostly sparse deciduous trees, while those in Vaihingen have denser foliage and canopy, as shown in Fig. 5. As a result, Trees and Low Vegetation in Potsdam share similar spectral characteristics with Low Vegetation in Vaihingen.
Finally, Fig. 10 illustrates some segmentation examples obtained in Task $2$ by various methods. It is clearly observed that the “Source-only” method suffered from the worst performance. Furthermore, despite the fact that existing UDA methods showed improvements of different degrees, they incurred deficiencies in the aspects of providing clear boundaries and differentiating visually similar categories. By leveraging its powerful global-local feature extraction capability and frequency decomposition frequency technology, the proposed FD-GLGAN achieved segmentation performance close to that by the supervised non-DA method.
Task 3 (V-IRRG to P-RGB): Task $3$ suffers from a domain shift problem similar to Task $2$, except that its source domain has a smaller data size than the target domain. Tabel 6 summarizes the results in Task $3$. From Table 6, it is observed that the proposed FD-GLGAN achieved the highest OA of $75.54\%$, mF1 of $76.27\%$, and mIoU of $62.17\%$, which stands for an improvement of $18.24\%$, $18.84\%$, and $20.27\%$ as compared to Advent, respectively. Furthermore, the segmentation accuracy on Tree and Low vegetation in Task $3$ has been substantially improved as compared to Task $2$. This is because these two categories in Vaihingen show much different spectral characteristics. Some segmentation samples are presented in Fig. 11 for visual inspection. In sharp contrast to other UDA methods, the proposed methods provided the most accurate semantic information with the smoothest boundaries.
In summary, the extensive experiments demonstrated that the proposed GLGAN and FD-GLGAN have the capability of overcoming domain shifts, leading to significantly improved semantic segmentation performance on remote sensing images.
5.6 Ablation Studies
Considering that we have proposed GLGAN based on GLTB, and HLFD based on decomposition technology, we conducted three sets of ablation studies to illustrate the necessity of each designation.
1) Global-Local information: To verify the effectiveness of the proposed backbone GLGAN, eight ablation experiments were performed as shown in Table 7. The baseline shown in the first row was obtained with a purely convolutional model composed of an encoder (Resnet101), a decoder (ASPP) and a CNN-based discriminator whereas the second row shows the case for the baseline with Resnet101 being replaced by the Swin Transformer for feature extraction. Furthermore, the third row shows the case in which the ASPP in the baseline was replaced by the GLTB-based decoder to recover the semantic features and predict the segmentation maps, while the CNN-based discriminator in the baseline was substituted with GLDis in the fourth row. The next three rows, i.e. the fifth row to the seventh row, are pairwise combinations of the Swin Transformer-based encoder, the GLTB-based decoder and GLDis. Inspection of the first four rows in Table 7 suggests that each proposed component is of great necessity in the proposed GLGAN.
Furthermore, the ablation experiments shown in the second and fifth rows confirmed the importance of the decoder in learning a transferable output space collocated with the Swin Transformer. Comparison of the results reported in the second and sixth rows, the third and seventh rows, and the fifth and last rows suggest that GLDis was able to improve overall segmentation performance, which provides strong evidence to support our motivation to strengthen the discriminator design under the GAN framework. Clearly, GLDis is a better choice than the conventional CNN-based discriminator. Table 7 confirmed that the proposed GLGAN is more efficient in image feature extraction and semantic information recovery by effectively exploiting intra- and inter-domain features.
Table 7: Ablation studies on the structure of GLGAN (%). Bold values are the best
Generator GLDis OA mF1 mIoU
encoder decoder
77.63 72.51 58.68
80.42 73.83 60.32
77.61 74.15 60.23
77.46 73.71 59.71
80.51 77.27 64.03
81.63 74.40 61.28
78.23 75.17 61.26
81.97 78.01 65.04
Table 8: Ablation studies for high/low-frequency of FD-GLGAN (%). HF denotes high-frequency and LF denotes low-frequency. Bold values are the best
HF LF OA mF1 mIoU
81.97 78.01 65.04
83.40 79.63 67.21
83.38 79.79 67.45
83.66 80.30 68.09
2) High/Low-frequency information: To verify the effectiveness of the two branches in HLFD, four ablation experiments were performed as shown in Table 8. The baseline shown in the first row was the results provided by GLGAN. The second row shows the case that only the high-frequency extractor is employed to align the high-frequency information whereas the third row shows the case that only the low-frequency extractor is employed to align the low-frequency information. The last row presents the results achieved by utilizing the complete HLFD. Inspection of the results suggests that in UDA tasks, the separate alignment of high- and low- frequency information is very important when learning cross-domain transferable features. We can also draw an inference that besides sharing similarities in global features, the source and target domains also exhibit resemblances in local features, such as the shape of tree crowns and vehicle outlines. These local characteristics of ground objects can be analyzed from a frequency perspective, thereby offering valuable insights for unlabeled target domains.
3) Multiscale features: Eight ablation experiments were performed as shown in Table 9 to verify the necessity of aligning multiscale features. The baseline shown in the first row was still the results provided by GLGAN. According to the experimental results we obtained, we present the ablation experiments in the order shown in Table 9. Specifically, we investigated the impact of a single scale, ranging from shallow features (scale 1) to deep features (scale 4), before progressively incorporating additional scales. There are two interesting conclusions confirmed by Table 9. Firstly, the alignment of multiscale features is necessary. Secondly, deep features hold greater significance compared to shallow features that is also widely recognized in other works (Chen et al., 2022; Ma et al., 2023b; Liu et al., 2022).
Table 9: Ablation studies on the multiscale features of FD-GLGAN (%). Bold values are the best
scale 1 scale 2 scale 3 scale 4 OA mF1 mIoU
81.97 78.01 65.04
82.89 78.82 66.13
83.06 78.83 66.26
82.44 79.29 66.68
83.34 79.45 66.96
83.22 79.83 67.35
83.39 80.17 67.92
83.66 80.30 68.09
In summary, the proposed GLGAN and FD-GLGAN achieved great segmentation performance as compared to other state-of-the-art UDA methods. Both quantitative results and visual inspection confirmed that GLTB-based backbone GLGAN and the frequency decomposition technology can largely bridge the segmentation performance gap between UDA methods and supervised methods.
Figure 12: Parameter analysis for high/low-frequency of FD-GLGAN.
5.7 Parameter Analysis
The sensitivity of hyper-parameter $\lambda^{f}$ is analyzed, as presented in Fig. 12. This parameter is designed to regulate the contribution derived from frequency decomposition. Inspection of the results in Fig. 12 indicates that a smaller value of $\lambda^{f}$ diminishes the importance of aligning high/low-frequency information, whereas a larger value may exaggerate the significance of the frequency alignment. Consequently, setting $\lambda^{f}\geq 0.05$ results in notable performance degradation. Conversely, performance demonstrates less sensitivity to $\lambda^{f}$ within the range $\lambda^{f}\in\left[0.008,0.02\right]$. Therefore, we set $\lambda^{f}=0.01$ in our experiments.
Figure 13: Visualization of feature maps before and after decomposition. The first row are the input images from the source and the target domains. The second to the fifth row corresponds to the four scales. Source domain feature maps: (a) before decomposition, (b) low-frequency and (c) high-frequency. Target domain feature maps: (d) before decomposition, (e) low-frequency and (f) high-frequency.
5.8 Visualization of Feature Decomposition
In this section, we will investigate the characteristics of the proposed multiscale HLFD modules by visually inspecting the heatmaps before and after frequency decomposition. Specifically, multiscale feature maps are first resized to a uniform image size, followed by averaging across the channel dimension. Subsequently, pseudo-color processing is applied. The results are depicted in Fig. 13. Upon inspecting the feature maps before and after decomposition in both the source and target domains, distinctions become apparent between low- and high-frequency information: low-frequency features possess global characteristics, whereas high-frequency features appear more discrete and localized. In particular, even the fourth scale (high-level) feature maps show significant differences. It fully demonstrates that our HLFD modules successfully decompose high- and low-frequency information.
Comparing feature maps between the source and target domains reveals that high-frequency features cross-domain share great similarities as well as the low-frequency features. It suggests that the model’s capacity to decompose features acquired from the labeled source domain can be successfully applied to the target domain. In other words, the model has learned domain-invariant features from both high- and low- frequency, which correspond to the similarity of spatial details across domains that we emphasized in Sec. 1 and the global context similarity that the existing works mainly explores. Our work has proved that both of these information are necessary for cross-domain alignment in the UDA method.
5.9 Model Complexity Analysis
Table 10 shows the computational complexity of all UDA methods discussed. Compared with the baseline Advent, the proposed general backbone GLGAN (Swin-Base) achieved greatly improved segmentation performance with significantly reduced computational complexity as its hierarchical window-based attention has linear computation complexity to input image size (Liu et al., 2021). More specifically, the computational complexity of GLGAN (Swin-Base) is only half that of Advent. Furthermore, GLGAN (Swin-Large) has computational complexity comparable to Advent. However, GLGAN (Swin-Large) requires much more memory to store its parameters. Overall, the proposed GLGAN harvests better performance at the cost of moderately increased computational complexity. The model complexity of FD-GLGAN is between GLGAN (Swin-Base) and GLGAN (Swin-Large), yet it exhibits superior performance. It proved that the HLFD module is of great significance in practical application. Furthermore, although we didn’t conduct FD-GLGAN (Swin-Large) limited by the experimental hardware, its performance holds promise and warrants further exploration.
Table 10: Comparison on computational complexity measured by a $256\times 256$ input on a single NVIDIA GeForce RTX 3090 GPU. Bold values are the best
Model Backbone Complexity (G) Parameters (M) Memory (MB) Speed (FPS)
ResUNet++
ResUNet++
241.82
35.46
4365
25.10
AdasegNet
ResNet101
47.59
42.83
2491
43.31
Advent
ResNet101
47.95
43.16
2491
30.65
GANAI
ResNet101
10.27
44.55
1075
6.30
TriADA
ResNet101
22.15
59.23
1075
4.54
CCAGAN
ResNet101
145.14
104.02
4503
13.04
MBATA-GAN
ResNet101
174.20
143.90
5295
8.11
GLGAN
Swin-Base
26.13
96.72
3639
12.50
Swin-Large
45.97
159.03
4840
9.95
FD-GLGAN
Swin-Base
34.26
138.68
3841
6.22
6 Conclusion
In this work, a novel plug-and-play technology named HLFD has been proposed to learn domain-invariant features through the decomposition of multiscale features from the frequency perspective. To further harness the potential of these decomposed features,a versatile UDA backbone, dubbed GLGAN, has been developed within the GAN framework. GLGAN is adept at extracting both intra- and inter-domain image features, leading to the generation of more precise segmentation maps. Additionally, a novel GLTB-based discriminator called GLDis has been established to enhance the generator’s capability to acquire domain-invariant representations. Extensive experiments on the well-known ISPRS Potsdam and Vaihingen datasets have confirmed that the proposed FD-GLGAN can substantially outperform existing UDA methods while greatly bridging the performance gap between the UDA methods and the supervised methods that utilize labeled data from the target domain. Finally, this work has successfully demonstrated the feasibility of solving the domain shift problem in the UDA tasks using highly efficient GLTB-based backbones and frequency decomposition technology, which will pave the way for future developments of more sophisticated UDA strategies.